# Standardizing the Measurement of Text Diversity: A Tool and Comparative Analysis

### v2

@@ 3.3 Use-Cases
Our implementation of compression ratios over parts-of-speech and tokens (along with BERTScore, Self-BLEU, and self-repetition) have been used in prior works (by other groups) to evaluate diversity in model evaluation and alignment (Lake et al., 2024; Moon et al., 2024; Fernandez et al., 2024), and for reporting diversity over synthetic datasets (Chang et al., 2024; Hastings et al., 2024). Due to its computational efficiency, compression ratios have also been used as optimization parameters in decoding strategies (Lanchantin et al., 2025).
Shaib et al. (2024) use the pattern analysis in diversity to measure and evaluate the prevalence of syntactic patterns in LLMs. Similarly, Wadhwa et al. (2025) extract part-of-speech patterns in distillation tasks for model attribution. Further, insights from our evaluation of diversity metrics have been used to inform how to report diversity with respect to text length and data sizes (Guo et al., 2023; Hastings et al., 2024).
Overall, prior work has shown how NLP researchers can greatly benefit from open access to diversity as well as the evaluation conducted across the implemented metrics. By releasing a user interface, we broaden use-cases for users who would like to explore patterns and diversity in text datasets without needing to use Python.
@@ 4.1 Data and Models
We compute diversity scores for the outputs of six instruction tuned models on the CNN/DailyMail (Hermann et al., 2015) and XSUM (Narayan et al., 2018) English news summarization datasets: Llama-2 (Touvron et al., 2023a), GPT-4 (OpenAI, 2023), FlanT5-XXL (Longpre et al., 2023), StableLM (Taori et al., 2023; Chiang et al., 2023; Anand et al., 2023), Mistral (Jiang et al., 2023), and StableBeluga (Touvron et al., 2023b; Mukherjee et al., 2023). [^8] All models—except GPT-4—downloaded from HuggingFace (https://huggingface.co/models). We selected these models to cover a range of availability (open and closed), and architectures (encoder-decoder, decoder-only). [^9] We use prompts for summarization provided by each model, where available. See Appendix C.4. The lengths of texts vary considerably by source, for reference and model-produced text alike, so we also note average lengths when reporting diversity.
Figure 3: Correlations between text diversity scores on CNN/DM. Compression ratio correlates strongly with most other diversity metrics.
@@ 5 Text Length as a Confounder
One would expect that human-authored texts would be more diverse than those produced by LLMs (with the caveat that the texts were scraped from the web, and so may contain HTML and page layout artefacts which might be repetitive (Salkar et al., 2022)). The human texts differ by length and the sources of longer texts appear to be less diverse. The association between the length of the produced texts and their diversity is similarly pronounced in the XSUM dataset, as seen Table 3. Text length as a confounder for diversity has been reported in prior work (Salkar et al., 2022), along with potential methods to adjust for this, e.g., sampling blocks of fixed size (Covington and McFall, 2010).
@@
Table 5 reports correlations between the number of words produced by each model and diversity scores. All scores of the token/type ratio family are highly correlated with length, while the pairwise similarity ones are only moderately correlated. Self-BLEU has low correlation with length.
@@
CR
CR: POS
NGD
Self-
Rep.
Hom.
(R-L)
Hom.
Self-
BLEU
MATTR HD-D
0.867 0.832 0.81 0.904 0.875 0.579 0.235 0.79 0.855
Table 5: Correlation between scores and total word counts (concatenated text) for CNN/Daily Mail.
6 Diversity of Model Summaries
The confound of length complicates reporting. On both CNN/DM and XSUM (cf. Tables 2 and 3), StableLM produces the longest summaries. All scores indicate that these are the least diverse, most likely due to the length confound. In both sets of results, we look for models that produce shorter summaries that are less diverse. These findings are notable and hold, despite length differences.
Three types of differences are marked in the tables. Model summaries that are shorter but less diverse than human summaries are marked in bold. Human texts here are written by journalists, so the expectation is that they would be more diverse. More bold entries in a column indicate that the score captures differences between human and machine diversity, a desirable trait. Underlined entries denote models that are less diverse than other models that produce longer summaries. The more underlined entries there are for a model, the more indicators there are that its output is less diverse. Asterisks mark models that appear more diverse than a human text of shorter length.
The most interesting diversity scores are those that capture differences between human and automatically produced text, without necessarily committing to an interpretation of which source is preferable. On the CNN/DM dataset, Hom. (BERT) and MATTR are the two scores that detect no differences between human and model texts. Compression ratio for part of speech sequences is the score that identifies the most differences between human and model-generated text. Self-repetition stands out as the only score that identifies model generated text as more diverse on the CNN/DM dataset. From this analysis, CR:POS and self-repetition emerge as prime candidates of reportable scores, while Hom. BERT is less useful.
7 Correlation Analysis
Compression ratio is highly to moderately correlated with other $n$-gram scores. The only weak correlations are with Self-BLEU and Hom. (BERT). Given the degenerate behavior of Hom. (BERT) on the analysis of summaries, reporting Self-BLEU only is advisable. Finally, self-repetition is only moderately correlated with with other scores, so it is informative to report this as a standard diversity score. Correlations are similar on XSUM summaries (Appendix 6), reinforcing the recommendations here. Diversity analysis on the CNN/DM and XSUM datasets did not indicate consistent system behavior. We report the analysis in Table 10.
Our guiding assumption is that output diversity and self-repetition are aspects of model behavior that are not captured by existing evaluation approaches. Here we directly test this assumption. We compute the system level correlation between the diversity scores and the traditional BERTScore and ROUGE evals, shown in Figure 4. Reference-based evaluations are only weakly correlated with the diversity metrics. Self-BLEU, however, is moderately anti-correlated with with both ROUGE-1 and BERTScore.
8 Truncating to Control Length
Model
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Article 2.268 5.25 2.763 0.676 0
Article (Lead 3) 2.274 5.25 2.762 0.658 0
Reference 2.189 5.179 2.763 0.674 0
Llama-2 2.96 5.627 2.847 0.674 0.001
GPT-4 2.287 5.376 2.761 0.672 0
FlanT5 2.288 5.389 2.779 0.673 0
StableLM 2.393 5.537 2.884 0.672 0.001
Mistral 2.32 5.415 2.812 0.67 0
StableBeluga 2.288 5.46 2.766 0.671 0
For each input for summarization, we truncate all summaries to the length of the shortest one produced by any of the sources as a crude means to remove the influence of length on scores. The resulting scores are directly comparable across sources, listed in Tables 6 and 4. Compression ratio and Self-BLEU scores indicate that model-produced text is less diverse than human text. Hom. (BERT) scores barely vary across sources, further supporting the recommendation that this is not a useful score to report. On the CNN/DM dataset, Self-BLEU indicates that Llama-2 and StableLM are the most repetitive models. Compression ratio also ranks these two models as the least diverse. The results are consistent on XSUM, but for that dataset Flan-T5 is also highly ranked and the most repetitive.
The truncation approach to control for length is not practical for published research or leaderboards. Introducing a new source of texts would require recomputing the scores for other sources one may want to compare with, which is impractical and sometimes impossible when the outputs from other sources are not available. Future research will have to search for more practical alternatives.
9 Discussion and Recommendations
The diversity package and platform provides a useful way to analyze and visualize diversity in datasets. For reporting metrics, our in-depth analyses reveal that compression ratio is an excellent score to report, easy to compute and strongly correlated with other scores used in past work. Compression ratio of part of speech sequences capture differences between human and model-generated text. Self-repetition zeros in only on repetition of longer $n$-grams across generations, and is only moderately correlated with compression ratios. Finally Self-BLEU is only weakly correlated with the previous three, so is a good complement score to report. In our analyses, we identified several drawbacks of BERTScore: it does not show differences between human and model-generated text and barely varies when adjusted for length.
Length of the analyzed text has to be reported alongside all these scores. When length differs, scores are not meaningfully comparable. Truncating and downsampling text is one way to produce a set of results that are intuitively comparable. Different random draws of the sample chosen to represent a dataset will likely differ in diversity; the selection may lead to unwarranted conclusions. Truncating texts prevents any possibility of discovering repetitive behavior towards the end of longer text. Future research into a principled solution for this problem is urgently needed. Despite all this, we were able to glean meaningful insights about differences in diversity between human and model-produced text for summaries, essays and instructions.
Finally, diversity offers a platform and package in which researchers from a variety of domains can use to facilitate evaluation.
10 Limitations
In the work presented here we do not explore human approaches to evaluating the diversity of collections of text. These are straightforward when the produced text is fairly short, as in judging the diversity of a set of questions generated for a given document (Sultan et al., 2020) or the diversity of possible continuation of a conversation (Tevet and Berant, 2021). Longer texts, as in the case of summaries, and larger collections, as in the case of instruction datasets are harder to judge for diversity.
Acknowledgements
We gratefully acknowledge the National Science Foundation (RI 2211954) for supporting this work.
Appendix A Appendix
Appendix B Run-Time Considerations
When analyzing the diversity of large volumes of text, run-time considerations become relevant. Figure 5 provides insights about the feasibility of obtaining scores for large samples [^10] Run on a single NVIDIA Quadro RTX 8000 GPU.. The compression ratio scores are fast, with text compression utilities specifically optimized for speed. Self-repetition takes longer but acceptable time. Self-BLEU and Hom. (BERT) are prohibitively slow.
Appendix C Broader Applications
The guiding motivation for this work has been to develop standardized and informed approach to the analysis the diversity of text produced by LLMs. The standardization of scores will facilitate analysis in broader settings. Here we provide two examples: human writing, with and without facilitation from a LLM, and instruction tuning datasets.
Human Story Writing
Padmakumar et al. (2023) presented an analysis of human-written stories, where people wrote either by themselves or with the help of GPT-3 or GPT-3.5 Turbo. They find that using LLMs as writing partners leads to greater homogenization of the stories. As reported by Padmakumar et al. (2023), we find that all diversity scores agree that people writing independently produce the more diverse texts (cf. Table 7). Here, story length is not an issue because the average length of stories in each setting are comparable: 375 words for writing without help, 372 words when writing with GPT-3 and 370 when writing with GPT-3.5.
Dataset
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Solo 2.901 5.314 5.873 0.604 0.018
GPT-3 2.940 5.371 5.911 0.613 0.020
InstructGPT 3.064 5.462 5.966 0.631 0.022
Instruction-tuning Datasets
The quality and diversity of instructions are likely to result in more robust and capable systems (Sanh et al., 2022; Mishra et al., 2022). We analyze the diversity of five instruction-tuning datasets: Open Assistant (Köpf et al., 2024), Super-NaturalInstructions (Wang et al., 2022), Unnatural Instructions (Honovich et al., 2023), Alpaca (Wang et al., 2023b), and Dolly (Conover et al., 2023).  [^11] We provide further details in Appendix C.6.
In Table 8 we report diversity scores. Here datasets are ordered by size; we therefore expect that scores will be sorted in diminishing order in each column. Only deviations from this ordering are reportable. We provide details about the number of instructions and words in Appendix C.6. Open Assistant instructions are remarkably diverse compared to the other datasets, and all diversity scores for it are more favorable than that for other datasets. Unnatural instructions are remarkable in the opposite direction, with outlier scores that are so much higher, they are likely not due to length entirely. We provide an analysis of the diversity scores with the length controlled in Appendix C.7.
@@ Appendix C Broader Applications
Dataset CR (↓) CR: POS (↓)
Self-Rep. (↓)
Open Assistant 2.886 6.731 3.969
Unnatural Instructions 4.191 7.278 9.868
Alpaca 3.119 6.61 3.105
Super-NaturalInstructions 2.675 5.749 3.456
Dolly 2.578 6.214 2.935
Table 8: Diversity scores for instruction datasets. We do not include Self-BLEU nor Hom. (BERT) due to long run times. Datasets are ordered by size and differ vastly in length, so only scores for which a smaller dataset is less diverse are meaningfully interpretable.
C.1 Human Assessments of Diversity
C.2 Examples of Repetitive Patterns
Table 9 show more examples of repeated sentence structures (using part-of-speech tags) from Padmakumar et al. (2023).
Dataset Token Repetition Text Pattern-Matched Text
GPT-3 "In my opinion…" 41/100 PRP VBZ RB JJ TO VB 15/100
"It is also vital to discern…"
"It is very easy to construct…"
"It is largely inappropriate to try…"
"It is morally acceptable to focus…"
PRP VBP IN DT NN IN 12/100
"I don’t like the damsel in…"
"I fear that a cycle of…"
"I feel that an acknowledgement of…"
"I find that the inflection of…"
Instruct-GPT "In my opinion…" 25/100 MD VB DT JJ NN IN 20/100
"It is important to…" 20/100 "…can have a huge variety of…"
"Up with the news…" 15/100 "…can have a negative effect on…"
"…can have a positive impact on…"
"…can have a sturdy framework for…"
PRP VBZ RB JJ TO VB 12/100
"It is also important to realize…"
"It is fairly common to hear…"
"It is indeed surprising to hear…"
"It is probably impossible to keep…"
"…it becomes very cringy to watch…"
Solo "In my opinion…" 22/100 PRP VBZ JJ TO VB IN 9/100
"In my opinion, the…" 13/100 "It is crucial to recognize that…"
"When it comes to…" 11/100 "It is crucial to remember that…"
"In my opinion, I…" 10/100 "It is unjustifiable to assume that…"
"It is important to acknowledge that…"
PRP VBP IN DT JJ NN 10/100
"I believe for the right person…"
"I do on a regular basis."
"I fall into the second group."
"I live in a small town…"
Table 9: Examples of exact text-match and repeating part-of-speech patterns in essays from Padmakumar and He (2023b). The number of times the pattern occurs is shown in parenthesis.
C.3 Correlation Between Metrics
@@
CR
CR: POS
NGD
Self-
Rep.
Hom.
(R-L)
Hom.
Self-
BLEU
MATTR HD-D
0.83 0.695 0.885 0.87 0.841 0.921 0.991 0.799 0.654
Table 10: Score correlations for each text diversity score between the CNN/DM and XSUM datasets.
C.4 Summarization Prompts
Table 11 details the prompts and format used to generate the summaries for the news datasets. We follow the formats recommended provided by each model, and insert the along with the instruction for summarization.
Model Model Size Prompt
Llama-2 7B [TEXT] [INST] Summarize the above text. [/INST]
GPT-4 - [TEXT]. Summarize the above text.
Flan-T5 11B Summarize this article: [TEXT]
StableLM 7B [TEXT] <|<|< |USER|>|>| >Summarize the above text.
<|<|< |ASSISTANT|>|>| >
Mistral 7B ### Instruction: Summarize the following: ### Input: [TEXT].
### Response:
Table 11: Prompts used for each model to generate a summary. [TEXT] is replaced with the input article.
C.5 XSUM Metrics
Figure 6: Correlation table between scores on XSUM.
C.6 Instruction Datasets, Details
Open Assistant
is a collection of crowdsourced instructions (Köpf et al., 2024). The data was collected under detailed guidelines and includes questions that reflect real-life situations.
Super-NaturalInstructions
A corpus comprising crowdsourced instructions that transform 200 benchmarks and intermediate evaluation results into a set of instructions and demonstrations (Wang et al., 2022).
Unnatural Instructions
An (almost) automatically created dataset, using instructions from the SuperNatural-Instructions dataset to automatically generate new instructions (Honovich et al., 2023). To increase diversity, each instruction was also paraphrased. Honovich et al. (2023) compare the diversity of instructions in Unnatural and Super-Natural Instructions with pairwise BERTScore similarities (within each dataset), and find that the similarities are much higher in Super-NaturalInstructions.
Alpaca
This dataset is created following the Self-instruct dataset (Wang et al., 2023b). GPT-3 was prompted to create instructions and demonstrations based on a seed of 175 human-written instructions. Crucially, the collection method includes a diversity filter, only including model-written instructions if their ROUGE-L similarity is less than 0.7 with an existing instruction. Length of instructions and demonstrations is also controlled for as a criterion for inclusion in the final instruction dataset.
Dolly
A set of human instructions and demonstrations, collected by Datrabricks employees (Conover et al., 2023). By design, they cover only eight classes of popular tasks: creative writing, closed and open QA, summarization, information extraction, classification and brainstorming.
Table 12 shows the number of instructions, the typical length of an instruction and average number of words per instruction set. All vary, making it even harder to control for length. Truncating makes less sense here, and down-sampling the number per instructions is counter-productive given our goal to understand the diversity of the entire dataset. We do make use of these instruments given the lack of alternatives, but note that more meaningful solutions are urgently needed.
Dataset # Instructions Avg. # Words Total # Words
Open Assistant 84,437 78.10 6,594,646
Unnatural Instructions 66,010 38.05 2,511,737
Alpaca 52,002 10.06 523,329
Super-NaturalInstructions 4550 92.58 421,228
Dolly 15,011 12.37 185,816
Table 12: Average number of words, and size of the instruction datasets. Numbers correspond to the training set available from Huggingface. For Super-NaturalInstructions, we filter for English-only instructions using the langdetect library.
C.7 Instruction Datasets, Length Controlled

### v3

@@ 3.3 Use-Cases
Our implementation of compression ratios over PoS tags and tokens (along with BERTScore, Self-BLEU, and self-repetition) have already been used in prior works (by other groups) to evaluate diversity in model evaluation and alignment (Lake et al., 2024; Moon et al., 2024; Fernandez et al., 2024), and for reporting diversity over synthetic datasets (Chang et al., 2024; Hastings et al., 2024). Due to its computational efficiency, compression ratios have also been used as optimization parameters in decoding strategies (Lanchantin et al., 2025).
shaib-etal-2024-detection use the pattern analysis in diversity to measure and evaluate the prevalence of syntactic patterns in LLMs. Wadhwa et al. (2025) extract PoS patterns in distillation tasks for model attribution. Further, insights from our evaluation of diversity metrics have informed how to report diversity with respect to text length and data sizes (Guo et al., 2024; Hastings et al., 2024).
@@ 4.1 Data and Models
We compute diversity scores for the outputs of six instruction tuned models on the CNN/DailyMail (Hermann et al., 2015) and XSUM (Narayan et al., 2018) English news summarization datasets: Llama-2 (Touvron et al., 2023a), GPT-4 (OpenAI, 2023), FlanT5-XXL (Longpre et al., 2023), StableLM (Taori et al., 2023; Chiang et al., 2023; Anand et al., 2023), Mistral (Jiang et al., 2023), and StableBeluga (Touvron et al., 2023b; Mukherjee et al., 2023). [^7] All models—except GPT-4—downloaded from HuggingFace (https://huggingface.co/models). We selected these models to cover a range of availability (open and closed), and architectures (encoder-decoder, decoder-only). The lengths of texts vary considerably by source, for reference and model-produced text alike, so we also note average lengths when reporting diversity.
Figure 2: Correlations between diversity scores on CNN/DM. CR correlates strongly with most other metrics.
@@ 5 Text Length as a Confounder
One would expect that human-authored texts would be more diverse than those produced by LLMs (with the caveat that the texts were scraped from the web, and so may contain HTML and page layout artifacts which might be repetitive (Salkar et al., 2022)). The human texts differ by length and the sources of longer texts appear to be less diverse.
Text length as a confounder for diversity has been reported in prior work (Salkar et al., 2022), along with methods to account for this, e.g., sampling blocks of fixed size (Covington and McFall, 2010).
@@
All scores of the token/type ratio family are highly correlated with length, while the pairwise similarity ones are only moderately correlated. Self-BLEU has low correlation with length.
6 Diversity of Model Summaries
The confound of length complicates reporting. On CNN/DM (cf. Table 1) StableLM produces the longest summaries. All scores indicate that these are the least diverse, probably due to length. Three types of differences are marked in the tables. Model summaries that are shorter but less diverse than human summaries are marked in bold. Human texts here are written by journalists, so the expectation is that they would be more diverse. More bold entries in a column indicate that the score captures differences between human and machine diversity, a desirable trait. Underlined entries denote models that are less diverse than other models that produce longer summaries. The more underlined entries there are for a model, the more indicators there are that its output is less diverse. Asterisks mark models that appear more diverse than a human text of shorter length.
The most interesting diversity scores are those that capture differences between human and automatically produced text. On the CNN/DM dataset, Hom. (BERT) and MATTR are the two scores that detect no differences between human and model texts. Compression ratio for part of speech sequences is the score that identifies the most differences between human and model-generated text. Self-repetition stands out as the only score that identifies model generated text as more diverse on the CNN/DM dataset. From this analysis, CR:POS and self-repetition emerge as prime candidates of reportable scores, while Hom. BERT is less useful.
7 Correlation Analysis
Compression ratio is highly to moderately correlated with other $n$-gram scores. The only weak correlations are with Self-BLEU and Hom. (BERT). Given the degenerate behavior of Hom. (BERT) on the analysis of summaries, reporting Self-BLEU only is advisable. Finally, self-repetition is only moderately correlated with other scores, and is therefore informative to report. Correlations are similar on XSUM summaries (Appendix 5).
8 Truncating to Control for Length
We truncate all summaries to the length of the shortest one produced by any source as a crude means to control scores for length. The resulting scores are directly comparable (see Table 5 in the Appendix). CRs and Self-BLEU scores indicate that model-generated text is less diverse than human text. Hom. (BERT) scores barely vary across sources. On the CNN/DM dataset, Self-BLEU indicates that Llama-2 and StableLM are the most repetitive models. CR also ranks these two models as the least diverse. The results are consistent on XSUM, but for that dataset Flan-T5 is also highly ranked and the most repetitive.
Truncation to control for length is impractical for published research or leaderboards. Introducing a new source of texts would require recomputing scores for other sources for comparison, which is sometimes impossible (when outputs from other sources are not available). Future research might search for more practical alternatives.
9 Discussion and Recommendations
The diversity package and platform provides a useful way to analyze and visualize diversity in datasets. Our analyses of metrics reveal that compression ratio (CR) is an excellent score to report, easy to compute and strongly correlated with other scores used in past work. CR of PoS sequences captures differences between human and model-generated text. Self-repetition focusses on repetition of longer $n$-grams across generations, and is only moderately correlated with CRs. Finally Self-BLEU is only weakly correlated with the previous three, so is a good complement score to report. We found BERTScore limited: It does not show differences between human and model-generated text and barely varies when adjusted for length.
Length of the analyzed text has to be reported alongside all these scores. When length differs, scores are not meaningfully comparable. Truncating text is one way to control for this. Different random draws of the sample chosen to represent a dataset may differ in diversity, in turn leading to unwarranted conclusions. Truncating texts prevents quantifying repetition towards the end of longer texts. Finally, diversity offers a platform and package in which researchers from a variety of domains can use to facilitate evaluation and gather insights about diversity between human- and model-produced texts.
10 Limitations
In this work, we do not attempt to measure human judgments of diversity, which are straightforward for short texts (e.g., questions) but far more difficult for longer summaries or large instruction datasets (tevet-berant-2021-evaluating); we leave this for future work. All evaluations are conducted in English.
Acknowledgements
We gratefully acknowledge the National Science Foundation (RI 2211954) for supporting this work. We thank Ramya Namuduri for contributing the QUDSim framework to the repository, and extend our appreciation to all other contributors.
Appendix
Appendix A Run-Time Considerations
Figure 3 provides insights about the feasibility of obtaining scores for large samples. [^8] Run on a single NVIDIA Quadro RTX 8000 GPU. The compression ratio scores are fast compared to other diversity measures.
Appendix B Correlations with Evaluation Metrics
Output diversity and self-repetition are aspects of model behavior that are not captured by existing evaluation approaches. We compute the system level correlation between the diversity scores and the traditional BERTScore and ROUGE evaluations, shown in Figure 4.
Appendix C Controlling for Length
Scores on CNN/DM summaries truncated to the shortest summary reveal a different model order with respect to diversity (Table 2).
Model
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Article 2.268 5.25 2.763 0.676 0
Article (Lead 3) 2.274 5.25 2.762 0.658 0
Reference 2.189 5.179 2.763 0.674 0
Llama-2 2.96 5.627 2.847 0.674 0.001
GPT-4 2.287 5.376 2.761 0.672 0
FlanT5 2.288 5.389 2.779 0.673 0
StableLM 2.393 5.537 2.884 0.672 0.001
Mistral 2.32 5.415 2.812 0.67 0
StableBeluga 2.288 5.46 2.766 0.671 0
Appendix D Additional Evaluations
Story Writing
Dataset CR (↓) CR: POS (↓)
Self-Rep. (↓)
Open Assistant 2.886 6.731 3.969
Unnatural Instructions 4.191 7.278 9.868
Alpaca 3.119 6.61 3.105
Super-NaturalInstructions 2.675 5.749 3.456
Dolly 2.578 6.214 2.935
Table 3: Diversity scores for instruction datasets. We do not include Self-BLEU nor Hom. (BERT) due to long run times.
padmakumar-etal-2023-investigating presented an analysis of human-written stories, where people wrote either by themselves or with the help of GPT-3 or GPT-3.5 Turbo. We also find that all diversity scores agree that people writing independently produce the more diverse texts (cf. Table 5). Length is not an issue because the average length of stories in each setting are comparable: 375 words for writing without help, 372 words when writing with GPT-3 and 370 when writing with GPT-3.5.
@@ Appendix D Additional Evaluations
Dataset
CR
CR: POS
Self-
Rep. (↓)
Hom.
(BERT) (↓)
Self-
BLEU (↓)
Solo 2.901 5.314 5.873 0.604 0.018
GPT-3 2.940 5.371 5.911 0.613 0.020
InstructGPT 3.064 5.462 5.966 0.631 0.022
Figure 5: Correlation table between scores on XSUM.
Instruction-tuning Datasets
The quality and diversity of instructions are likely to result in more robust and capable systems (Sanh et al., 2022; mishra-etal-2022-cross). We analyze the diversity of five instruction-tuning datasets: Open Assistant (Kopf et al., 2023), Super-NaturalInstructions (wang-etal-2022-super), Unnatural Instructions (honovich-etal-2023-unnatural), Alpaca (wang-etal-2023-self-instruct), and Dolly (Conover et al., 2023) (Table 3).
Open Assistant instructions are remarkably diverse compared to the other datasets across all diversity scores. Unnatural instructions are remarkable in the opposite direction. We provide an analysis of the diversity scores with the length controlled in Appendix D.2.
@@
D.1 Correlation Between Metrics
@@
D.2 Instruction Datasets, Length Controlled